OpenStudy (anonymous):

2 cards are randomly drawn frm a deck of 52 cards.find the probability that both cards will be greater then 3 but less than 8?

OpenStudy (anonymous):

how many cards fit the bill? 4,5,6,7 of each suit, so 16 in total. the probability first card fits this restriction is \[\frac{16}{52}\] and the probabilty the second card does GIVEN THE FIRST CARD DOES

OpenStudy (anonymous):

is \[\frac{15}{51}\]

OpenStudy (anonymous):

answer is \[\frac{16}{52}\times \frac{15}{51}\]

OpenStudy (anonymous):

which certainly reduces if you like

OpenStudy (anonymous):

it is without replacement?

OpenStudy (anonymous):

drawing two cards is without replacement right? first card, then second card

OpenStudy (anonymous):

ok.ok

OpenStudy (anonymous):

any good site forpracticing probability?